Using only R -Code

#The rain probability for the month of July in the city of Portland is 72%

#(a) What is the probability that it rains exactly 11 days

#(b) What is the probability that it rains 11 or fewer days?

#(c) What is the probability that it rains 27 or more days?

#(d)Generate 10000 random numbers using rnorm function. Show how this distribution would look like. Add in the plot the mean, median and 20% trimmed mean (provide these values).
Answer #1

#a
dbinom(11,31,0.72)

#b
pbinom(11,31,0.72)

#c
1-pbinom(26,31,0.72)

#d
z=rnorm(10000)
plot(z)
M1=mean(z)
Med=median(z)
M2=mean(z,trim=0.2) # 20% trim mean
M1
Med
M2
X=c(M1,Med,M2)

points(X,pch=24,col="red")

Output

[1] 2.00244e-05
[1] 2.495013e-05
[1] 0.04033098
[1] mean= -0.0007766617
[1] median= 0.002720538
[1] trimmed mean= 0.002691458